# Adjoint log canonical foliated singularities on surfaces

### v1

@@ 4.4.4. Summary of the classification
Fix $\epsilon\in(0,\tfrac{1}{3})$. Let $(Y,\mathcal{G},p)$ be a germ of a foliated surface, and assume that $p$ is an $\epsilon$-adjoint log canonical singularity of $(Y,\mathcal{G})$. Let
@@
be the minimal resolution (cf. Definition 2.5) of $(Y,\mathcal{G},p)$, with exceptional divisor $E=\bigcup_{i=1}^{r}E_{i}$. Then $E$ is one of the following:
  1. (1)
    An $\mathcal{F}$-chain. (See Figure 3)
  2. (2)
    A chain of (2)(-2)-$\mathcal{F}$-curves. (See Definition 2.18.)
  3. (3)
    $E=\Theta_{1}+\Theta_{2}+C$, where $C$ is a smooth rational $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$, meeting transversely two $\mathcal{F}$-chains $\Theta_{1}$ and $\Theta_{2}$. Moreover, $\Theta_{1}=\Gamma_{11}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{11}$ with self-intersection 2-2, and $\Theta_{2}$ is one of:
@@
    2. (b)
    3. (c)
  4. (4)
    $E=C+\Theta_{1}+\Theta_{2}+\Xi_{1}$ is a single fork centered at $C$, where $C$ is a smooth rational $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$, $\Theta_{1}$ and $\Theta_{2}$ are $\mathcal{F}$-chains, and $\Xi_{1}$ is a chain of (2)(-2)-$\mathcal{F}$-curves. Moreover, $\Theta_{1}=\Gamma_{11}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{11}$ with self-intersection 2-2, and $\Theta_{2}$ is one of:
    1. (a)
      $\Theta_{2}=\Gamma_{21}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{21}$ with self-intersection 2-2; (see Figure 5)
    2. (b)
    3. (c)
  5. (5)
  6. (6)
    An $\mathcal{F}$-star chain, that is, a chain with exactly one non-$\mathcal{F}$-invariant curve $E_{k}$, for some $k\in\{1,\dots,r\}$. Moreover, ${\rm tang}(\mathcal{F},E_{k})=0$, and $E_{1}+\cdots+E_{k-1}$ (resp. $E_{k+1}+\cdots+E_{r}$) is an $\mathcal{F}$-chain whose first curve is $E_{k-1}$ (resp. $E_{k+1}$). (See Figure 8.)
  7. (7)
  8. (8)
    An $\mathcal{F}$-star graph of type $(l;m_{1},\dots,m_{l})$ centered at $C$ (cf. Definition 2.22), where $C$ is a smooth rational non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C)=0$, and $(l;m_{1},\dots,m_{l})$ is one of:
    1. (a)
@@
    (See Definition 2.22 and Figure 7.)
Moreover, the following hold:
@@
    Type (5), (7), (8-b), and (8-c) are not $\epsilon$-adjoint klt.
@@
    Type (3-b), (3-c), (4-b) and (4-c) cannot occur if $\epsilon\in(0,1/5)$.
@@
The statement follows by combining Proposition 4.8, Proposition 4.9, Proposition 4.10, and Proposition 4.13. ∎
5. $\epsilon$-adjoint canonical singularities with $\epsilon\in(0,1/4)$
Throughout this section, we fix $\epsilon\in(0,1/4)$. We characterize those configurations in Theorem 4.14 that are $\epsilon$-adjoint canonical.
@@ 5. $\epsilon$-adjoint canonical singularities with $\epsilon\in(0,1/4)$
Type (5), (7), (8-b), and (8-c) are not $\epsilon$-adjoint canonical.
@@
Type (3-b), (3-c), (4-b), and (4-c) are not $\epsilon$-adjoint canonical.
@@
Suppose that $p$ is an $\epsilon$-adjoint canonical singularity. Then $E$ is a $(K^{\prime})^{\geq 0}$-graph and hence also a $W^{\geq 0}$-graph. Clearly, $W\cdot\Gamma\geq 0$ for any component $\Gamma\subset E$ with $\Gamma\neq C$. We claim that $W\cdot C>0$, which yields a contradiction.
@@
  • If $\Theta_{2}$ is of type (3-b) or (4-b), then
@@
  • If $\Theta_{2}$ is of type (3-c) or (4-c), then
@@
Type (8-a) is not $\epsilon$-adjoint canonical.
@@
We follow the notation of Lemma 5.2. It suffices to show $W\cdot C>0$, where $C$ is as in type (8) of Theorem 4.14 and
$W:=K^{\prime}-\sum_{i=1}^{3}M(K^{\prime},\Theta_{i}).$
In case (8-a), we have $e_{C}:=-C^{2}\geq 2$, since the intersection matrix of $C+\Theta_{1}+\Theta_{2}+\Theta_{3}$ is negative definite. Thus,
@@
Case $(2,2,n)$, $n\geq 2$. We have
$M(K^{\prime},\Theta_{1})=\frac{1}{2}\Gamma_{11},\qquad M(K^{\prime},\Theta_{2})=\frac{1}{2}\Gamma_{21}.$
Thus,
$M(K^{\prime},\Theta_{1})\cdot C=M(K^{\prime},\Theta_{2})\cdot C=\frac{1}{2}.$
On the other hand, since the coefficients of $M(K^{\prime},\Theta_{3})$ are at most (n1)/n(n-1)/n, we have
@@
Case $(2,3,n)$, $n=3,4,5$. Here $\Theta_{1}=\Gamma_{11}$ with $\Gamma_{11}^{2}=-2$, and $\Theta_{2}$ is either
@@
$M(K^{\prime},\Theta_{1})C=\frac{1}{2},\quad M(K^{\prime},\Theta_{2})C=\frac{1-\epsilon}{3},\quad M(K^{\prime},\Theta_{3})C\leq\frac{n-1}{n}\leq\frac{4}{5},$
which imply
@@
$M(K^{\prime},\Theta_{1})C=\frac{1}{2},\quad M(K^{\prime},\Theta_{2})C=\frac{2}{3},\quad M(K^{\prime},\Theta_{3})C\leq\frac{n-1}{n}\leq\frac{4}{5},$
which imply
@@
If type (3-a) (resp. type (4-a)) is $\epsilon$-adjoint canonical, then every component of $E$ has self-intersection 2-2.
@@
The proof is similar to that of Lemma 5.2. ∎
@@
If type (6) is $\epsilon$-adjoint canonical, then the non-$\mathcal{F}$-invariant curve $E_{k}$ is a (1)(-1)-curve.
@@
We use the notation as in Lemma 5.2. Let $C:=E_{k}$. In this case, $C$ meets $\mathcal{F}$-chains $\Theta_{1},\dots,\Theta_{\ell}$ with $\ell\leq 2$. Set
$W:=K^{\prime}-\sum_{i=1}^{\ell}M(K^{\prime},\Theta_{i}).$
Assume $C^{2}\leq-2$. Then
@@
Since $M(K^{\prime},\Theta_{i})\cdot C<1$ for all $i$, we obtain
$W\cdot C>2-\ell\geq 0,$
contradicting the fact that $E$ is a $W^{\geq 0}$-graph. ∎
The above lemmas exclude most configurations in Theorem 4.14 under the $\epsilon$-adjoint canonical condition for $\epsilon\in(0,1/4)$. The remaining cases yield the following classification.
@@
be the minimal resolution, with exceptional divisor $E=\bigcup_{i=1}^{r}E_{i}$. Then $E$ is one of the following:
@@
    An $\mathcal{F}$-dihedral graph, all of whose components have self-intersection 2-2. (See Figure 5.)
@@
    An $\mathcal{F}$-star chain, that is, a chain with exactly one smooth rational non-$\mathcal{F}$-invariant curve $E_{k}$ with ${\rm tang}(\mathcal{F},E_{k})=0$, such that the two sides are $\mathcal{F}$-chains. Moreover, in this case, $E_{k}^{2}=-1$. (See Figure 8.)
Here, types (1)(4) are canonical foliated singularities, and types (3) and (4) are not $\epsilon$-adjoint terminal.
Appendix A Log canonical surface singularities
In this appendix, we give a proof of the classification of log canonical surface singularities, following an approach analogous to the program developed in Section 4.
Let $Y$ be a normal surface and let $p$ be a log canonical singularity of $Y$. Let
$\pi:X\to Y$
be a minimal resolution of $(Y,p)$ with exceptional curves $\mathfrak{E}=\bigcup_{i\in I}E_{i}$, and set
$E:=\sum_{i\in I}E_{i}.$
Then we can write
$K:=K_{X}+\sum_{i\in I}E_{i}=\pi^{*}K_{Y}+\sum_{i\in I}a_{i}E_{i},$
so that $E$ is a $K^{\geq 0}$-graph.
A.1. $K$-non-positive curves in $E$
Lemma A.1.
Let $C\subset E$ be an irreducible $\mathcal{F}$-invariant curve. Then
(A.1) $KC=2p_{a}(C)-2+d(C).$
Moreover, $KC\leq 0$ if and only if $C$ is one of:
  1. (1)
    $d(C)=0$:
    1. (a)
      a smooth rational curve. In this case, KC=2KC=-2.
    2. (b)
      a smooth elliptic curve. In this case, $KC=0$.
    3. (c)
      a rational curve with a single node. In this case, $KC=0$.
  2. (2)
    $C$ is a smooth rational curve with $d(C)=1$. In this case, KC=1KC=-1.
  3. (3)
    $C$ is a smooth rational curve with $d(C)=2$. In this case, $KC=0$.
A.2. Special $K^{>0}$-chain
Proposition A.2.
Suppose $E$ has at least two components. Let $F=C_{1}+\cdots+C_{r}$ be a special $K^{>0}$-chain with the first curve $C_{1}$. Then $F$ is a Hirzebruch-Jung chain with
$d(C_{1})=1,\qquad d(C_{i})=2,\quad i\geq 2,$
and
$KC_{1}=1,\qquad KC_{2}=0,\quad i\geq 2.$

### v2

@@ 4.4.4. Summary of the classification
Combining Propositions 4.9, 4.10, 4.11, and 4.14, we obtain a complete classification of admissible configurations.
@@
Fix $\epsilon\in(0,1/3)$. Let $(Y,\mathcal{G},p)$ be a germ of a foliated surface such that $p$ is $\epsilon$-adjoint log canonical. Let
@@
be the minimal resolution, and write $E=\bigcup_{i=1}^{r}E_{i}$ for the exceptional divisor.
Then the exceptional divisor $E$ is one of the following configurations.
  1. (I)
    Foliated canonical configurations.
    • (1)
      A chain of smooth rational $\mathcal{F}$-invariant curves, consisting of:
      • (a)
        A $\mathcal{F}$-chain (Figure 1).
      • (b)
        Two (1)(-1)-$\mathcal{F}$-curves of self-intersection 2-2 joined by a bad tail (Figure 2).
      • (c)
        A chain of (2)(-2)-$\mathcal{F}$-curves.(See Definition 2.16)
    • (2)
      An $\mathcal{F}$-dihedral fork (Figure 3).
    • (3)
  2. (II)
    Foliated lc (non-canonical) configurations.
    • (1)
      An $\mathcal{F}$-star chain centered at a smooth rational curve: a chain $E=E_{1}+\cdots+E_{r}$ of smooth rational curves such that for some index $k$, the curve $E_{k}$ is non-$\mathcal{F}$-invariant with ${\rm tang}(\mathcal{F},E_{k})=0$, and the two subchains $E_{1}+\cdots+E_{k-1}$ and $E_{k+1}+\cdots+E_{r}$ (if nonempty) are $\mathcal{F}$-chains whose first curves are $E_{k-1}$ and $E_{k+1}$ respectively.
      (See Figure 6.)
    • (2)
      An $\mathcal{F}$-star fork of type $(l;m_{1},\dots,m_{l})$ centered at a smooth rational curve $C$ (cf. Definition 2.22), where $C$ is non-$\mathcal{F}$-invariant and ${\rm tang}(\mathcal{F},C)=0$, and where $(l;m_{1},\dots,m_{l})$ is one of the following:
@@
        $l=3$ and $(m_{1},m_{2},m_{3})=(2,2,n)$ for $n\geq 2$, or $(2,3,3)$, $(2,3,4)$, $(2,3,5)$;
@@
      (See Figure 5.)
    • (3)
  3. (III)
    Boundary configurations.
    • (1)
      $E=\Theta_{1}+\Theta_{2}+C$, where $C$ is a smooth rational $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$, meeting transversely two $\mathcal{F}$-chains $\Theta_{1}$ and $\Theta_{2}$ of type $(2,3)$. Moreover, $\Theta_{1}=\Gamma_{11}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{11}$ with self-intersection 2-2, and $\Theta_{2}$ is one of:
      1. (a)
      2. (b)
    • (2)
      $E=C+\Theta_{1}+\Theta_{2}+\Xi_{1}$ is a single fork centered at $C$, where $C$ is a smooth rational $\mathcal{F}$-invariant curve with ${\rm Z}(\mathcal{F},C)=3$, $\Theta_{1}$ and $\Theta_{2}$ are $\mathcal{F}$-chains of type $(2,3)$, and $\Xi_{1}$ is a chain of (2)(-2)-$\mathcal{F}$-curves. Moreover, $\Theta_{1}=\Gamma_{11}$ for some (1)(-1)-$\mathcal{F}$-curve $\Gamma_{11}$ with self-intersection 2-2, and $\Theta_{2}$ is one of:
      1. (a)
      2. (b)
In addition, the following properties hold.
@@
    Type (I) is foliated canonical, Type (II) is foliated lc (non-canonical), and Type (III) is not foliated lc. In particular, Type (III) cannot occur if $\epsilon<1/5$.
@@
    Among Type (I) and (II),
    • (I-1), (I-2), (II-1) and (II-2-a) are $\epsilon$-adjoint klt;
    • (I-3), (II-2-b),(II-2-c) and (II-3) are not $\epsilon$-adjoint klt.
@@
All possible configurations of the exceptional divisor $E$ are covered by Proposition 4.9, Proposition 4.10, Proposition 4.11, and Proposition 4.14. Hence, by collecting the conclusions of these propositions, we obtain both the classification of $E$ and the additional properties listed in the theorem. ∎
Corollary 4.16.
Fix $\epsilon\in(0,1/5)$. Suppose $p$ is an $\epsilon$-adjoint log canonical singularity of $(X,\mathcal{F})$. Then $p$ is log canonical both as a foliated singularity and as a surface singularity. Moreover, if $p$ is $\epsilon$-adjoint klt, then $p$ is klt as a singularity of $X$.
Proof.
It follows directly from Theorem 4.15, together with the classification of lc surface singularities (Theorem 2.24). ∎
Remark 4.17.
The classification carried out in this section should be understood as a finite numerical reduction procedure on the negative definite exceptional divisor $E$, rather than as a mere enumeration of combinatorial possibilities.
From this viewpoint, the same reduction mechanism, with minor adjustments to the numerical parameters, recovers the classical classifications of log canonical surface singularities (cf. Theorem 2.24) and log canonical foliated surface singularities (cf. Theorem 2.25), which fit naturally into the same numerical reduction framework.
5. $\epsilon$-adjoint canonical singularities for $\epsilon\in(0,1/4)$
Throughout this section, we fix $\epsilon\in(0,1/4)$. Under the $\epsilon$-adjoint canonical condition, we determine which of the configurations listed in Theorem 4.15 are ruled out and which can still occur as $\epsilon$-adjoint canonical singularities.
@@ 5. $\epsilon$-adjoint canonical singularities for $\epsilon\in(0,1/4)$
Types (I-3), (II-2-b), (II-2-c), and (II-3) are not $\epsilon$-adjoint canonical.
@@
Types (III) is not $\epsilon$-adjoint canonical.
@@
Suppose that $p$ is an $\epsilon$-adjoint canonical singularity. Then $E$ is a $(K^{\prime})^{\geq 0}$-graph and hence also a $W^{\geq 0}$-graph. Clearly, $W\cdot\Gamma\geq 0$ for any component $\Gamma\subset E$ with $\Gamma\neq C$. We claim that $W\cdot C>0$, contradicting the assumption that $E$ is a $W^{\geq 0}$-graph.
@@
If $\Theta_{2}$ is of type (III-1-a) or (III-2-a), then
@@
If $\Theta_{2}$ is of type (III-1-b) or (III-2-b), then
@@
Type (II-2-a) is not $\epsilon$-adjoint canonical.
@@
We follow the notation of Lemma 5.2. Let
$W:=K^{\prime}-\sum_{i=1}^{3}M(K^{\prime},\Theta_{i}),$
where $\Theta_{1},\Theta_{2},\Theta_{3}$ are the $\mathcal{F}$-chains attached to $C$ as in type (II-2-a) of Theorem 4.15. Similar to the proof of Lemma 5.2, it suffices to show that $W\cdot C>0$.
In case (II-2-a), we have $e_{C}:=-C^{2}\geq 2$, since the intersection matrix of $C+\Theta_{1}+\Theta_{2}+\Theta_{3}$ is negative definite. Hence
@@
Case $(2,2,n)$, $n\geq 2$.
We have
$M(K^{\prime},\Theta_{1})=\tfrac{1}{2}\Gamma_{11},\qquad M(K^{\prime},\Theta_{2})=\tfrac{1}{2}\Gamma_{21},$
so
$M(K^{\prime},\Theta_{1})\cdot C=M(K^{\prime},\Theta_{2})\cdot C=\tfrac{1}{2}.$
Since the coefficients of $M(K^{\prime},\Theta_{3})$ are at most (n1)/n(n-1)/n, we obtain
@@
Case $(2,3,n)$, $n=3,4,5$.
Here $\Theta_{1}=\Gamma_{11}$ with $\Gamma_{11}^{2}=-2$, and $\Theta_{2}$ is either
@@
$M(K^{\prime},\Theta_{1})\cdot C=\tfrac{1}{2},\quad M(K^{\prime},\Theta_{2})\cdot C=\frac{1-\epsilon}{3},\quad M(K^{\prime},\Theta_{3})\cdot C\leq\frac{n-1}{n}\leq\frac{4}{5},$
hence
@@
$M(K^{\prime},\Theta_{1})\cdot C=\tfrac{1}{2},\quad M(K^{\prime},\Theta_{2})\cdot C=\tfrac{2}{3},\quad M(K^{\prime},\Theta_{3})\cdot C\leq\frac{n-1}{n}\leq\frac{4}{5},$
hence
@@
Let $E=C_{1}+\cdots+C_{r}$ be as in type (I-1-c) of Theorem 4.15. If $p=\pi(E)$ is $\epsilon$-adjoint canonical, then
$(-C_{1}^{2},\ldots,-C_{r}^{2})=(2,\ldots,2)\quad\text{or}\quad(2,\ldots,2,3,2,\ldots,2,1).$
If moreover $p=\pi(E)$ is $\epsilon$-adjoint terminal, then $p$ is a smooth point of $Y$, and around $p$ the foliation $\mathcal{G}$ is locally defined (cf. [Bru15, pp. 8]) by
$\omega=(ry+x^{r})\,{\rm d}x-x\,{\rm d}y.$
@@
Note that each $C_{i}$ is not $\mathcal{F}$-exceptional, and $E$ is a $(K^{\prime})^{\geq 0}$-graph for $K^{\prime}=K_{\mathcal{F}}+\epsilon K_{X}$.
If $C_{i}^{2}\leq-2$ for all $i$, then
$K^{\prime}C=K_{\mathcal{F}}C-\epsilon K_{X}C=\epsilon K_{X}C\geq 0$
and equality holds if and only if $C_{i}^{2}=-2$. Thus, the fact that $E$ is a $(K^{\prime})^{\geq 0}$-graph implies $C_{i}^{2}=-2$ for all $i$. In this case, $E$ is a $(K^{\prime})^{=0}$-graph.
Suppose that $C_{i}^{2}=-1$ for some $i$. Then necessarily $i=1$ or $i=r$; otherwise, $C_{i}$ would be $\mathcal{F}$-exceptional, a contradiction. By the separatrix theorem (cf. Theorem 2.10) and Lemma 2.18, there exists at most one such curve $C_{i}$. Without loss of generality, assume $C_{r}^{2}=-1$. By Lemma 2.18, there exists exactly one singularity of $\mathcal{F}$ on $C_{r}$, which is a saddle-node whose weak separatrix is contained in $C_{r}$.
For $i\leq r-1$, we have $C_{i}^{2}\leq-2$. Let $l$ be the integer such that $C_{i}^{2}=-2$ for $l\leq i\leq r-1$ and $C_{l-1}^{2}\leq-3$. Contracting $C_{l}+C_{l+1}+\cdots+C_{r}$, we obtain a birational morphism
$\sigma\colon(X,\mathcal{F})\longrightarrow(X^{\prime},\mathcal{F}^{\prime}).$
Let
$E^{\prime}=C^{\prime}_{1}+\cdots+C^{\prime}_{l-1},\qquad C^{\prime}_{i}:=\sigma_{*}(C_{i}).$
Then $E^{\prime}$ is a $(K_{\mathcal{F}^{\prime}}+\epsilon K_{X^{\prime}})^{\geq 0}$-graph. Moreover, $(C^{\prime}_{i})^{2}\leq-2$ and $K_{\mathcal{F}^{\prime}}\cdot C^{\prime}_{i}=0$ for all $i\leq l-1$. It follows that $(C^{\prime}_{i})^{2}=-2$ for all $i\leq l-1$. Therefore,
$(-C_{1}^{2},\ldots,-C_{r}^{2})=(2,\ldots,2,3,2,\ldots,2,1).$
If $l\geq 2$, then $p=\pi(E)$ is not $\epsilon$-adjoint terminal. If $l=1$, then $p$ is a smooth point of $Y$. In this case, around $p$ the foliation $\mathcal{G}$ is locally defined (cf. [Bru15, pp. 8]) by
$\omega=(ry+x^{r})\,{\rm d}x-x\,{\rm d}y.$
@@
If type (I-1-b) (resp. type (I-2)) is $\epsilon$-adjoint canonical, then every component of $E$ has self-intersection 2-2.
@@
Let $C$ denote the tail in type (I-1-b) (resp. the fork center in type (I-2)). Let $\Theta_{1},\Theta_{2}$ be the $\mathcal{F}$-chains attached to $C$ as in type (I-1-b) (resp. (I-2)). Then $\Theta_{i}=\Gamma_{i}$, where $\Gamma_{i}$ is a (1)(-1)-$\mathcal{F}$-curve with self-intersection 2-2 for $i=1,2$. Moreover, every component of $E$ has self-intersection $\leq-2$ in both types (I-1-b) and (I-2).
We follow the notation of Lemma 5.2. Set
$W:=K^{\prime}-\sum_{i=1}^{2}M(K^{\prime},\Theta_{i})=K^{\prime}-\frac{1}{2}\Gamma_{1}-\frac{1}{2}\Gamma_{2}.$
Since $E$ is a $(K^{\prime})^{\geq 0}$-graph, it is also a $W^{\geq 0}$-graph.
We compute
$W\Gamma_{1}=W\Gamma_{2}=0$
and
$WC=K^{\prime}C-1=K_{\mathcal{F}}C+\epsilon K_{X}C-1=\epsilon(-2-C^{2})\geq 0.$
For any remaining curve $C^{\prime}$ in type (I-2), we have
$WC^{\prime}=K^{\prime}C^{\prime}=K_{\mathcal{F}}C^{\prime}+\epsilon K_{X}C^{\prime}=\epsilon(-2-(C^{\prime})^{2})\geq 0.$
Thus, by Lemma 2.26, the fact that $E$ is a $W^{\geq 0}$-graph implies $C^{2}=-2$ and $(C^{\prime})^{2}=-2$. Therefore, every component of $E$ has self-intersection 2-2, as claimed. ∎
Lemma 5.6.
Let $E=E_{1}+\cdots+E_{k}+\cdots+E_{r}$ be as in type (II-1) of Theorem 4.15, where $E_{k}$ is the unique non-$\mathcal{F}$-invariant curve. If $p=\pi(E)$ is $\epsilon$-adjoint canonical, then $E_{k}^{2}=-1$ and $r\geq 2$.
Proof.
We use the notation as in Lemma 5.2. Let $C:=E_{k}$. In this setting, the curve $C$ meets $l\leq 2$ $\mathcal{F}$-chains, denoted by $\Theta_{1},\dots,\Theta_{l}$. For each $i$, let $\Theta^{\prime}_{i}$ be a maximal subchain of $\Theta_{i}$ such that $\Theta^{\prime}_{i}$ is a $(K^{\prime})^{\geq 0}$-graph. Set
$W:=K^{\prime}-\sum_{i=1}^{l}M(K^{\prime},\Theta^{\prime}_{i}).$
Then $E$ is a $W^{\geq 0}$-graph.
By the definition of $M(K^{\prime},\Theta^{\prime}_{i})$ (cf. Definition 3.1) and the maximality of the $\Theta^{\prime}_{i}$’s, we have:
  • $W\cdot\Gamma=0$ for any irreducible curve $\Gamma\subset\bigcup_{i}\Theta^{\prime}_{i}$, and
  • $W\cdot\Gamma\geq 0$ for any irreducible curve $\Gamma\subset\bigl(\bigcup_{i}\Theta_{i}\bigr)\setminus\bigl(\bigcup_{i}\Theta^{\prime}_{i}\bigr)$.
Assume that $C^{2}\leq-2$. Then
@@
Since $M(K^{\prime},\Theta^{\prime}_{i})\cdot C<1$ for all $i$, it follows that
$W\cdot C>2-l\geq 0,$
which contradicts the fact that $E$ is a $W^{\geq 0}$-graph. Thus $C^{2}=-1$.
Next, we show that $l\geq 1$, or equivalently $r\geq 2$. If $l=0$, then $W=K^{\prime}$. In this case,
$W\cdot C=K^{\prime}\cdot C=1-\epsilon>0,$
which again contradicts the fact that $E$ is a $W^{\geq 0}$-graph. Therefore, $l\geq 1$, and hence $r\geq 2$. ∎
The above lemmas exclude all configurations in Theorem 4.15 except the following five types. We therefore obtain the complete classification of $\epsilon$-adjoint canonical singularities for $\epsilon\in(0,1/4)$.
@@
be the minimal resolution of $(Y,\mathcal{G},p)$, with exceptional divisor $E=\bigcup_{i=1}^{r}E_{i}$. Then $E$ is one of the following:
@@
    An $\mathcal{F}$-dihedral fork, all of whose components have self-intersection 2-2. (See Figure 3.)
@@
    An $\mathcal{F}$-star chain centered at a smooth rational curve, that is, a chain with exactly one smooth rational non-$\mathcal{F}$-invariant curve $E_{k}$ with ${\rm tang}(\mathcal{F},E_{k})=0$, such that the two sides are $\mathcal{F}$-chains. Moreover, in this case, $E_{k}^{2}=-1$ and $r\geq 2$. (See Figure 6.)
Moreover, the following holds:
  • (i)
    Types (1)(4) are foliated canonical and type (5) is foliated log canonical.
  • (ii)
    In type (2), the point $p\in Y$ is either a Du Val singularity or a smooth point of $Y$. In the former case, $p$ is not $\epsilon$-adjoint terminal.
  • (iii)
    Types (3) and (4) are not $\epsilon$-adjoint terminal; moreover, they are Du Val singularities of $Y$.
Remark 5.8.
By Theorem 5.7, the instability of the canonical condition under adjoint perturbations is entirely due to configuration (5).
There is no uniform constant $\epsilon_{0}>0$ that excludes configuration (5) for all $\epsilon<\epsilon_{0}$ (see Example 6.1). However, for each fixed index $i(\mathcal{G})$, there exists a constant $\epsilon_{0}(i(\mathcal{G}))>0$ such that configuration (5) does not occur whenever $\epsilon<\epsilon_{0}(i(\mathcal{G}))$. For instance, [LWX25, Proposition 1.9] shows that if $K_{\mathcal{G}}$ is big, one may take $\epsilon_{0}(i(\mathcal{G}))=\frac{1}{3i(\mathcal{G})}$.
Corollary 5.9.
Fix $\epsilon\in(0,1/4)$. Suppose $p$ is an $\epsilon$-adjoint canonical singularity of $(X,\mathcal{F})$. Then $p$ is log canonical as a foliated singularity and klt as a surface singularity.